### v2

# Rigidity Results for Compact Submanifolds with Pinched Ricci Curvature in Euclidean and Spherical Space Forms

Abstract.
For compact submanifolds in Euclidean and Spherical space forms with Ricci curvature bounded below by a function $\alpha(n,k,H,c)$ of mean curvature, we prove that the submanifold is either isometric to the Einstein Clifford torus, or a topological sphere for the maximal bound $\alpha(n,[\frac{n}{2}],H,c)$, or has up to $k$-th homology groups vanishing. This gives an almost complete (except for the differentiable sphere theorem) characterization of compact submanifolds with pinched Ricci curvature, generalizing celebrated rigidity results obtained by Ejiri, Xu-Tian, Xu-Gu, Xu-Leng-Gu, Vlachos, Dajczer-Vlachos.
Key words and phrases:
Pinched Ricci curvature; rigidity theorem; Einstein.
2010 Mathematics Subject Classification:
53C20, 53C24, 53C40.
the corresponding author.
J. Q. Ge is partially supported by NSFC (No. 12171037) and the Fundamental Research Funds for the Central Universities.
Y. Zhou is partially supported by NSFC (No. 12171037, 12271040), China Postdoctoral Science Foundation (No. BX20230018) and National Key R&\&&D Program of China 2020YFA0712800.
1. Introduction
In 1979, Ejiri [6] obtained a celebrated rigidity theorem for $n(\geq 4)$-dimensional, simply connected, compact, orientable, minimal immersed submanifolds in the unit sphere with pinched Ricci curvature $\operatorname{Ric}\geq n-2$. In 2011, Xu-Tian [16] removed the assumption that the submanifold is simply connected. Later, Xu-Gu [14] generalized Ejiri’s rigidity theorem to $n(\geq 3)$-dimensional, compact, orientable submanifolds in a real space form $F^{n+m}(c)$ with $c+H^{2}>0$ under the assumption $\operatorname{Ric}\geq(n-2)\left(c+H^{2}\right)$ and parallel mean curvature vector field $\mathcal{H}$ ($H:=|\mathcal{H}|$). In the same paper, Xu-Gu raised a conjecture, indicating that the above generalization should also hold without the assumption of parallel mean curvature vector field. Recently, Dajczer-Vlachos [3] affirmed Xu-Gu’s conjecture for the case $n\geq 4$, expect for the part of diffeomorphic to spheres. The same philosophy was used in a series of rigidity results [1, 2, 4] for submanifolds with pinched Ricci curvature in different settings. For the case of spherical space forms [1, 2], the core techniques are different improvements of the following theorem:
Theorem 1.1.
(((([12, Theorem 2])))) Let $M^{n}$ be an $n(\geq 4)$-dimensional compact submanifold of the unit sphere $\mathbf{S}^{n+m}$. Assume that the Ricci curvature satisfies
(1.1) $\operatorname{Ric}>b(n,k,H):=\frac{n(k-1)}{k}+\frac{n(k-1)H}{2k^{2}}\left(nH+\sqrt{n^{2}H^{2}+4k(n-k)}\right),$
where $k$ is an integer such that $2\leq k\leq\left[\frac{n}{2}\right]$. Then $\pi_{1}(M)=0$ and $H_{p}(M;\mathbb{Z})=H_{n-p}(M;\mathbb{Z})=0$ for all $1\leq p\leq k$.
According to [12], the function $b(n,k,H)$ comes from an expression for the Ricci curvature of the standard immersion of the torus $S^{k}(r)\times S^{n-k}(\sqrt{1-r^{2}})$ into the unit sphere $\mathbf{S}^{n+1}$, where $S^{k}\left(r\right)$ denotes the $k$-dimensional sphere of radius $r$. It is natural to compare $b(n,k,H)$ with another lower bound $\alpha(n,k,H,c)$ which was used in a similar homology vanishing result [15, Lemma 4.1] under Ricci curvature pinching condition. We then define and describe the function $\alpha(n,k,H,c)$.
Throughout this paper, we use the notation $F^{n}(c)$ to denote the simply connected, $n$-dimensional real space form with constant sectional curvature $c$. In addition, we set
$\alpha(n,k,H,c):=\bigg{(}n-1-\dfrac{k(n-k)(n-2)}{k(n-2k)+n(k-1)(n-k)}\bigg{)}\left(c+H^{2}\right).$
For the Einstein torus $S^{k}\left(\sqrt{\frac{k-1}{n-2}}\,r\right)\times S^{n-k}\left(\sqrt{\frac{n-k-1}{n-2}}\,r\right)\subset S^{n+1}\left(r\right)$ $(2\leq k\leq\left[\frac{n}{2}\right])$ embedded in $F^{n+m}(c)$ with mean curvature $H$, a direct calculation shows that its Ricci curvature can be expressed by $\operatorname{Ric}=\frac{n-2}{r^{2}}=\alpha(n,k,H,c)$.
Now we compare $\alpha(n,k,H,1)$ and $b(n,k,H)$ for $n\geq 4$ and $2\leq k\leq\left[\frac{n}{2}\right]$. By a direct calculation, we obtain
$b(n,k,H)-\alpha(n,k,H,1)=$ $\frac{n(k-1)}{2k^{2}}H\sqrt{n^{2}H^{2}+4k(n-k)}$
$+\frac{n(k-1)(n-2k)}{k\big{(}(n+2)\gamma(k)+2n\big{)}}\bigg{(}\frac{n}{2k}\gamma(k)H^{2}-(n-k)\bigg{)},$
where $\gamma(k):=k(n-k)-n\geq n-4\geq 0$. Hence we have the following results.
  • $(1)$
    $\alpha(n,k,0,1)\geq b(n,k,0)$ and the equality holds if and only if $n=2k$.
  • $(2)$
    $\alpha(n,k,H,1)<b(n,k,H)$ for sufficiently large $H$.
  • $(3)$
    If $n=2k$, then we have $\alpha(n,\frac{n}{2},H,1)\leq b(n,\frac{n}{2},H)$ and the equality holds if and only if $H=0$.
In the following, we state the main theorem of this paper which improves the pinched condition of [15, Lemma 4.1] and presents a complete rigidity result.
Theorem 1.2.
Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 5$, $c\geq 0$, be an isometric immersion of a compact, connected Riamannian manifold. If the Ricci curvature of $M$ satisfies
(*) $\operatorname{Ric}\geq\alpha(n,k,H,c)=\bigg{(}n-1-\dfrac{k(n-k)(n-2)}{k(n-2k)+n(k-1)(n-k)}\bigg{)}\left(c+H^{2}\right)$
for an integer $k$ with $2\leq k\leq\left[\frac{n}{2}\right]$, then one of the following holds:
  • $(1)$
    $\pi_{1}(M)=0$ and
    $H_{q}(M;\mathbb{Z})=H_{n-q}(M;\mathbb{Z})=0\ \mbox{for each}\ 1\leq q\leq k.$
    If in addition $k<\left[\frac{n}{2}\right]$, then $H_{n-k-1}(M;\mathbb{Z})$ is torsion-free.
  • $(2)$
    $M$ is isometric to the Einstein hypersurface $S^{k}\left(\sqrt{\frac{k-1}{n-2}}\,r\right)\times S^{n-k}\left(\sqrt{\frac{n-k-1}{n-2}}\,r\right)$ in the totally umbilical sphere $S^{n+1}\left(r\right)\subset F^{n+m}(c)$, where $r=\sqrt{\frac{n-2}{\alpha(n,k,H,c)}}$ with constant $H$. In this case, $\operatorname{Ric}=\alpha(n,k,H,c)$ on $M$.
The next corollary analyses the special case $k=\left[\frac{n}{2}\right]$ in Theorem 1.2. Thus we can give another proof for [3, Theorem 1] when $n\geq 5$. Different from Dajczer-Vlachos’ method, when $M$ is not a topological sphere, we show that $M$ has parallel mean curvature vector field by a careful study of the first normal bundle.
Corollary 1.3.
Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 5$, $c\geq 0$, be an isometric immersion of a compact, connected Riemannian manifold. If the Ricci curvature of $M$ satisfies
$\operatorname{Ric}\geq(n-2)\left(c+H^{2}\right),$
then $M$ is either homeomorphic to the $n$-sphere or isometric to the minimal Clifford hypersurface $S^{k}\left(r/\sqrt{2}\right)\times S^{k}\left(r/\sqrt{2}\right)$ in the totally umbilical sphere $S^{n+1}\left(r\right)\subset F^{n+m}(c)$, where $n=2k$ and $r=\frac{1}{\sqrt{c+H^{2}}}$.
For the case of odd-dimensional manifold, we can obtain the following rigidity result under a weaker pinched condition, which improves [15, Theorem 1.4]. Other rigidity results for odd-dimensional submanifolds in the unit sphere are also obtained in [1, 2, 11].
Corollary 1.4.
Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 5$, $c\geq 0$, be an isometric immersion of a compact, connected, odd-dimensional Riemannian manifold. If the Ricci curvature of $M$ satisfies
$\operatorname{Ric}\geq\left(n-2-\frac{4}{n^{3}-2n^{2}-n-2}\right)\left(c+H^{2}\right),$
then $M$ is either homeomorphic to the $n$-sphere or isometric to the Einstein hypersurface $S^{k}\left(\sqrt{\frac{k-1}{n-2}}\,r\right)\times S^{n-k}\left(\sqrt{\frac{n-k-1}{n-2}}\,r\right)$ in the totally umbilical sphere $S^{n+1}\left(r\right)\subset F^{n+m}(c)$, where $n=2k+1$ and $r=\sqrt{\frac{n-2}{\alpha(n,k,H,c)}}$.
The rest of this paper is organized as follows. In Section 2, we give some preliminaries and show the core lemma of this paper (Lemma 2.5). In Section 3, we prove Theorem 1.2, Corollary 1.3 and Corollary 1.4.
2. Preliminaries
Let $f:M^{n}\rightarrow F^{n+m}(c)$ be an isometric immersion, and let $\operatorname{II}:TM\times TM\rightarrow NM$ be the second fundamental form. For an arbitrary fixed point $x\in M$, let $\{e_{1},\cdots,e_{n}\}$ and $\{\xi_{1},\cdots,\xi_{m}\}$ be orthonormal bases of $T_{x}M$ and $N_{x}M$, respectively. Write
$\operatorname{II}(e_{i},e_{j})=\sum_{\alpha=1}^{m}h_{ij}^{\alpha}\xi_{\alpha}\ \mbox{for}\ 1\leq i,j\leq n.$
For each $1\leq\alpha\leq m$, let
$H_{\alpha}:=\left(h_{ij}^{\alpha}\right)_{n\times n}\ \mbox{and}\ c_{\alpha}:=\frac{1}{n}\operatorname{tr}H_{\alpha}=\frac{1}{n}\sum_{i=1}^{n}h_{ii}^{\alpha}.$
Then the mean curvature vector field is given by
(2.1) $\mathcal{H}=\frac{1}{n}\sum_{i=1}^{n}\operatorname{II}(e_{i},e_{i})=\frac{1}{n}\sum_{\alpha=1}^{m}\sum_{i=1}^{n}h_{ii}^{\alpha}\xi_{\alpha}=\sum_{\alpha=1}^{m}c_{\alpha}\xi_{\alpha}.$
Let $H=|\mathcal{H}|$ be the mean curvature, and let $S=|\operatorname{II}|^{2}$ be the squared length of the second fundamental form. Then the Gauss equation implies that
(2.2) $\operatorname{Ric}(e_{i},e_{j})=(n-1)c\delta_{ij}+\sum_{\alpha=1}^{m}\bigg{(}nc_{\alpha}h_{ij}^{\alpha}-\sum_{k=1}^{n}h_{ik}^{\alpha}h_{jk}^{\alpha}\bigg{)},$
(2.3) $\rho=n(n-1)c+n^{2}H^{2}-S,$
where $\operatorname{Ric}$ and $\rho$ are the Ricci tensor and the scalar curvature of $M$, respectively. In addition, we use the notation $\operatorname{Ric}(X):=\operatorname{Ric}(X,X)$ for $X\in TM$ with $|X|=1$. For each $x\in M$, the subspace
$N_{1}(x)=\operatorname{Span}\{\operatorname{II}(X,Y):X,Y\in T_{x}M\}$
is called the first normal space of $f$ at $x$. An isometric immersion $f$ is said to be $1$-regular if the first normal space $N_{1}(x)$ has constant dimension independent of $x\in M$. For a $1$-regular isometric immersion $f$, the first normal spaces form a subbundle $N_{1}$ of the normal bundle $NM$, called the first normal bundle of $f$.
Besides, we also need the following terminology. A normal vector $\eta$ at $x\in M$ is called a principal normal of $f$ at $x$ if the subspace
$E_{\eta}(x):=\left\{X\in T_{x}M:\operatorname{II}(X,Y)=\langle X,Y\rangle\eta\ \mbox{for each}\ Y\in T_{x}M\right\}$
is nontrivial. A smooth normal vector field $\eta$ on $M$ is called a principal normal vector field of $f$ with multiplicity $k>0$ if $E_{\eta}(x)$ has constant dimension $k$ (and hence is a smooth distribution). If in addition $\eta$ is parallel (with respect to the normal connection) along $E_{\eta}$, then it is called a Dupin principal normal vector field. According to [5, Proposition 1.22], every principal normal vector field with multiplicity $k\geq 2$ is a Dupin principal normal vector field.
Next, we will present several lemmas used in the proof of Theorem 1.2 and Corollary 1.3. The following Lemma 2.1 was first proved by Lawson-Simons [8] for the case $c>0$, and then by Xin [13] for the case $c=0$ when the inequality (##\##2.1) is strict on the whole submanifold. Elworthy-Rosenberg [7] observed that the result still holds by only requiring the inequality (##\##2.1) is strict at one point.
Lemma 2.1.
(((([7, 8, 13])))). Let $f:M^{n}\rightarrow F^{n+m}(c)$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that
(##\##) $\Theta_{q}:=\sum_{i=1}^{q}\sum_{j=q+1}^{n}\left(2|\operatorname{II}(e_{i},e_{j})|^{2}-\langle\operatorname{II}(e_{i},e_{i}),\operatorname{II}(e_{j},e_{j})\rangle\right)\leq q(n-q)c$
holds for any point $x\in M$ and any orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$, where $q$ is an integer such that $11\leq q\leq n-$ 1. If there exists a point of $M$ such that the inequality ((((#italic-#\##2.1)))) is strict, then $H_{q}(M;\mathbb{Z})=H_{n-q}(M;\mathbb{Z})=0$.
Using Lemma 2.1, Dajczer-Vlachos [3] improved [14, Lemma 2.2] to the following Lemma 2.2.
Lemma 2.2.
(((([3, Lemma 4])))) Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 4$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that the Ricci curvature of $M$ satisfies
(2.4) $\operatorname{Ric}\geq\frac{n(n-1)}{n+2}\left(c+H^{2}\right).$
[…]
The following Lemma 2.3 concerns the estimate of $\alpha(n,k,H,c)$, which will be used several times later. For $n\geq 5$, let $\varphi:[2,\frac{n}{2}]\rightarrow\mathbb{R}$ be the smooth function given by
$\varphi(s):=\frac{s(n-s)}{s(n-2s)+n(s-1)(n-s)}.$
Then we have
$\alpha(n,k,H,c)=\big{(}n-1-(n-2)\varphi(k)\big{)}\left(c+H^{2}\right),$
and we will only consider the case of $c+H^{2}>0$ in the following lemma.
Lemma 2.3.
The function $\varphi(s)$ is strictly monotone decreasing, and thus $\alpha(n,k,H,c)$ is strictly monotone increasing for $2\leq k\leq\left[\frac{n}{2}\right]$. In addition, we have
(2.5) $(n-3)\left(c+H^{2}\right)<\alpha(n,k,H,c)\leq(n-2)\left(c+H^{2}\right)$
for any $2\leq k\leq\left[\frac{n}{2}\right]$, and the equality holds if and only if $n=2k$.
Proof.
A direct calculation shows that
$\varphi^{\prime}(s)=\frac{-n^{2}(n-2s)}{(s(n-2s)+n(s-1)(n-s))^{2}}<0$
for $s\in[2,\frac{n}{2})$. It follows that $\varphi(s)$ is strictly monotone decreasing, and thus $\alpha(n,k,H,c)$ is strictly monotone increasing for $2\leq k\leq\left[\frac{n}{2}\right]$. In addition, we have
(2.6) $\alpha(n,k,H,c)\leq\big{(}n-1-(n-2)\varphi(n/2)\big{)}\left(c+H^{2}\right)=(n-2)\left(c+H^{2}\right),$
(2.7) $\alpha(n,k,H,c)$ $\geq\big{(}n-1-(n-2)\varphi(2)\big{)}\left(c+H^{2}\right)=\big{(}n-1-\frac{2(n-2)^{2}}{n^{2}-8}\big{)}\left(c+H^{2}\right)$
$>(n-3)\left(c+H^{2}\right),$
where the last inequality in (2.7) follows from the fact that $(n-2)^{2}<n^{2}-8$ for $n\geq 5$. It is easy to see that the equality holds in (2.6) if and only if $n=2k$. ∎
In the study of the first normal bundle $N_{1}$, we need the following Lemma 2.4 to show that $N_{1}$ is parallel.
Lemma 2.4.
(((([10, Lemma 5])))) Let $f:M^{n}\rightarrow F^{n+m}(c)$ be a $1$-regular immersion. If for every point of $M$, there exists an orthonormal frame $\{E_{1},\cdots,E_{n}\}$ such that for each $1\leq i\leq n$, $\operatorname{II}(E_{i},E_{j})=0$ for any $j\neq i$ and $\operatorname{II}(E_{i},E_{i})=\sum_{j\neq i}a_{j}\operatorname{II}(E_{j},E_{j})$ for some real numbers $a_{j}$’s, then the first normal bundle $N_{1}$ is parallel.
Finally, we state and prove the core lemma of this paper which is based on the proof of [15, Lemma 4.1].
Lemma 2.5.
Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 5$, be an isometric immersion satisfying
(*) $\operatorname{Ric}\geq\alpha(n,k,H,c)=\bigg{(}n-1-\dfrac{k(n-k)(n-2)}{k(n-2k)+n(k-1)(n-k)}\bigg{)}\left(c+H^{2}\right)$
at a point $x\in M$ for an integer $k$ such that $2\leq k\leq\left[\frac{n}{2}\right]$.
  • (1)
    The inequality ((((#italic-#\##2.1)))) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. Moreover, if the inequality ((((*2.5)))) is strict at $x$, then the inequality ((((#italic-#\##2.1)))) is strict for $q=k$ at $x$.
  • (2)
    Suppose that the equality holds in ((((#italic-#\##2.1)))) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$. Then for any orthonormal basis $\{\xi_{1},\cdots,\xi_{m}\}$ of $N_{x}M$, the corresponding matrix
    $H_{\alpha}=\operatorname{diag}(\lambda_{\alpha}I_{k},\mu_{\alpha}I_{n-k})\ \mbox{for each}\ 1\leq\alpha\leq m,$
    where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and $n-k$, respectively. Furthermore, if $c+H^{2}>0$ at $x$, then $\eta_{1}=\sum_{\alpha=1}^{m}\lambda_{\alpha}\xi_{\alpha}$ and $\eta_{2}=\sum_{\alpha=1}^{m}\mu_{\alpha}\xi_{\alpha}$ are distinct principal normals of $f$ at $x$ such that $E_{\eta_{1}}(x)=\operatorname{Span}\{e_{1},\cdots,e_{k}\}$ and $E_{\eta_{2}}(x)=\operatorname{Span}\{e_{k+1},\cdots,e_{n}\}$. In addition,
    $\operatorname{Ric}(e_{i},e_{j})=\delta_{ij}\alpha(n,k,H,c)\ \mbox{for any}\ 1\leq i,j\leq n.$
Proof.
To prove (1), we will estimate $\Theta_{q}$ for any $2\leq q\leq k$. Let $\{e_{1},\cdots,e_{n}\}$ and $\{\xi_{1},\cdots,\xi_{m}\}$ be orthonormal bases of $T_{x}M$ and $N_{x}M$, respectively. It follows from (2.1-2.2) that $\sum\limits_{\alpha=1}^{m}c_{\alpha}^{2}=H^{2}$ and
(2.8) $\operatorname{Ric}(e_{i})=(n-1)c+\sum_{\alpha=1}^{m}\bigg{(}nc_{\alpha}h^{\alpha}_{ii}-\sum_{j=1}^{n}(h^{\alpha}_{ij})^{2}\bigg{)}.$
On one hand, we have
(2.9) $\Theta_{q}$ $=2\sum_{\alpha=1}^{m}\sum_{j=q+1}^{n}\sum_{i=1}^{q}(h_{ij}^{\alpha})^{2}-\sum_{\alpha=1}^{m}\sum_{j=q+1}^{n}\sum_{i=1}^{q}h_{ii}^{\alpha}h_{jj}^{\alpha}$
$=\sum_{\alpha=1}^{m}\bigg{(}2\sum_{j=q+1}^{n}\sum_{i=1}^{q}(h_{ij}^{\alpha})^{2}-\big{(}\sum_{i=1}^{q}h_{ii}^{\alpha}\big{)}\big{(}nc_{\alpha}-\sum_{i=1}^{q}h_{ii}^{\alpha}\big{)}\bigg{)}$
$\leq\sum_{\alpha=1}^{m}\bigg{(}2\sum_{j=q+1}^{n}\sum_{i=1}^{q}(h_{ij}^{\alpha})^{2}-nc_{\alpha}\sum_{i=1}^{q}h_{ii}^{\alpha}+q\sum_{i=1}^{q}(h_{ii}^{\alpha})^{2}\bigg{)}$
$\leq q\sum_{i=1}^{q}\big{(}(n-1)c-\operatorname{Ric}(e_{i})\big{)}+n(q-1)\sum_{\alpha=1}^{m}\sum_{i=1}^{q}c_{\alpha}h_{ii}^{\alpha}$
$\leq q^{2}\big{(}(n-1)\left(c+H^{2}\right)-\operatorname{Ric}_{\min}(x)\big{)}$
$\quad\,-q(n-q)H^{2}+n(q-1)\sum_{\alpha=1}^{m}\sum_{i=1}^{q}c_{\alpha}(h_{ii}^{\alpha}-c_{\alpha}),$
where $\operatorname{Ric}_{\min}(x):=\min\left\{\operatorname{Ric}(X):X\in T_{x}M,|X|=1\right\}$. On the other hand, we have
(2.10) $\Theta_{q}$ $=\sum_{\alpha=1}^{m}\bigg{(}2\sum_{j=q+1}^{n}\sum_{i=1}^{q}(h_{ij}^{\alpha})^{2}-\frac{n-q}{n}\big{(}\sum_{i=1}^{q}h_{ii}^{\alpha}\big{)}\big{(}nc_{\alpha}-\sum_{i=1}^{q}h_{ii}^{\alpha}\big{)}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \ -\frac{q}{n}\big{(}\sum_{j=q+1}^{n}h_{jj}^{\alpha}\big{)}\big{(}nc_{\alpha}-\sum_{j=q+1}^{n}h_{jj}^{\alpha}\big{)}\bigg{)}$
$\leq\sum_{\alpha=1}^{m}\bigg{(}2\sum_{j=q+1}^{n}\sum_{i=1}^{q}(h_{ij}^{\alpha})^{2}-(n-q)c_{\alpha}\sum_{i=1}^{q}h_{ii}^{\alpha}+\frac{q(n-q)}{n}\sum_{i=1}^{q}(h_{ii}^{\alpha})^{2}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \ -qc_{\alpha}\sum_{j=q+1}^{n}h_{jj}^{\alpha}+\frac{q(n-q)}{n}\sum_{j=q+1}^{n}(h_{jj}^{\alpha})^{2}\bigg{)}$
$\leq\frac{q(n-q)}{n}S-\sum_{\alpha=1}^{m}\bigg{(}qnc_{\alpha}^{2}+(n-2q)c_{\alpha}\sum_{i=1}^{q}h_{ii}^{\alpha}\bigg{)}$
$\leq q(n-q)\bigg{(}(n-1)\left(c+H^{2}\right)-\operatorname{Ric}_{\min}(x)\bigg{)}$
$\quad\,-q(n-q)H^{2}-(n-2q)\sum_{\alpha=1}^{m}\sum_{i=1}^{q}c_{\alpha}(h_{ii}^{\alpha}-c_{\alpha}).$
From (2.9-2.10), the assumption (*2.5) and Lemma 2.3, we obtain
(2.11) $\Theta_{q}$ $\leq\frac{n-2q}{q(n-2)}\bigg{(}q^{2}\big{(}(n-1)\left(c+H^{2}\right)-\operatorname{Ric}_{\min}(x)\big{)}$
$\qquad\qquad\quad\ \ -q(n-q)H^{2}+n(q-1)\sum_{\alpha=1}^{m}\sum_{i=1}^{q}c_{\alpha}(h_{ii}^{\alpha}-c_{\alpha})\bigg{)}$
$\quad+\frac{n(q-1)}{q(n-2)}\bigg{(}q(n-q)\big{(}(n-1)\left(c+H^{2}\right)-\operatorname{Ric}_{\min}(x)\big{)}$
$\qquad\qquad\qquad\ \,-q(n-q)H^{2}-(n-2q)\sum_{\alpha=1}^{m}\sum_{i=1}^{q}c_{\alpha}(h_{ii}^{\alpha}-c_{\alpha})\bigg{)}$
$=\frac{q(n-2q)+n(q-1)(n-q)}{n-2}\big{(}(n-1)\left(c+H^{2}\right)-\operatorname{Ric}_{\min}(x)\big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\ \ -q(n-q)H^{2}$
$\leq\frac{\big{(}q(n-2q)+n(q-1)(n-q)\big{)}k(n-k)}{k(n-2k)+n(k-1)(n-k)}\left(c+H^{2}\right)-q(n-q)H^{2}$
$\leq\frac{\big{(}q(n-2q)+n(q-1)(n-q)\big{)}q(n-q)}{q(n-2q)+n(q-1)(n-q)}\left(c+H^{2}\right)-q(n-q)H^{2}$
$=q(n-q)c$
for any $2\leq q\leq k$. Note that the last inequality in (2.11) is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. Hence the inequality (##\##2.1) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$. If in addition the inequality (*2.5) is strict at $x$, then the second inequality in (2.11) is strict, and thus the inequality (##\##2.1) is strict for $q=k$ at $x$.
Next, we prove $(2)$. Assume that the equality holds in (##\##2.1) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$. Then all inequalities in (2.9-2.11) become equalities. The first inequality in (2.10) implies that for each $1\leq\alpha\leq m$, there exist $\lambda_{\alpha},\mu_{\alpha}\in\mathbb{R}$ such that
(2.12) $h_{ii}^{\alpha}=\lambda_{\alpha}\ \mbox{for any}\ 1\leq i\leq k,$
(2.13) $h_{jj}^{\alpha}=\mu_{\alpha}\ \mbox{for any}\ k+1\leq j\leq n,$
and the second inequality in (2.10) implies that for each $1\leq\alpha\leq m$,
(2.14) $\bigg{(}\frac{k(n-k)}{n}-1\bigg{)}h_{ij}^{\alpha}=0\ \mbox{for any}\ 1\leq i\leq k,\ k+1\leq j\leq n,$
(2.15) $h_{ii^{\prime}}^{\alpha}=h_{jj^{\prime}}^{\alpha}=0\ \mbox{for any}\ 1\leq i\neq i^{\prime}\leq k,\ k+1\leq j\neq j^{\prime}\leq n.$
Since $2\leq k\leq\left[\frac{n}{2}\right]$, we have $\frac{k(n-k)}{n}>1$ for $n\geq 5$, and thus it follows from (2.12-2.15) that
$H_{\alpha}=\operatorname{diag}(\lambda_{\alpha}I_{k},\mu_{\alpha}I_{n-k})\ \mbox{for each}\ 1\leq\alpha\leq m,$
Then by (2.2), we have
$\operatorname{Ric}(e_{i},e_{j})=0\ \mbox{for any}\ 1\leq i\neq j\leq n.$
In addition, combining the last inequality in (2.10) and the second inequality in (2.11), we obtain
(2.16) $\operatorname{Ric}(e_{i})=\alpha(n,k,H,c)\ \mbox{for each}\ 1\leq i\leq n.$
Furthermore, we consider the normal vectors $\eta_{1}=\sum_{\alpha=1}^{m}\lambda_{\alpha}\xi_{\alpha}$ and $\eta_{2}=\sum_{\alpha=1}^{m}\mu_{\alpha}\xi_{\alpha}$. If $\eta_{1}=\eta_{2}$, then by (2.2), we have $\operatorname{Ric}(e_{i})=(n-1)\left(c+H^{2}\right)$ for each $1\leq i\leq n$, which contradicts (2.16) and Lemma 2.3 if $c+H^{2}>0$ at $x$. Hence, $\eta_{1}\neq\eta_{2}$ and for any tangent vectors $X=\sum_{i=1}^{n}x_{i}e_{i},Y=\sum_{i=1}^{n}y_{i}e_{i}\in T_{x}M$, we have
$\operatorname{II}(X,Y)$ $=\sum_{\alpha=1}^{m}\sum_{i,j=1}^{n}x_{i}y_{j}h_{ij}^{\alpha}\xi^{\alpha}=\sum_{\alpha=1}^{m}\Big{(}\sum_{i=1}^{k}x_{i}y_{i}\lambda_{\alpha}\xi^{\alpha}+\sum_{j=k+1}^{n}x_{j}y_{j}\mu_{\alpha}\xi^{\alpha}\Big{)}$
$=\Big{(}\sum_{i=1}^{k}x_{i}y_{i}\Big{)}\eta_{1}+\Big{(}\sum_{j=k+1}^{n}x_{j}y_{j}\Big{)}\eta_{2}.$
It follows that $\eta_{1}$, $\eta_{2}$ are principal normals of $f$ at $x$ such that $E_{\eta_{1}}(x)=\operatorname{Span}\{e_{1},\cdots,e_{k}\}$ and $E_{\eta_{2}}(x)=\operatorname{Span}\{e_{k+1},\cdots,e_{n}\}$. ∎
3. Proof of Theorems
In this section, we will prove Theorem 1.2, Corollary 1.3 and Corollary 1.4.
3.1. Proof of Theorem 1.2
Proof.
[…]
(3.1) $\alpha(n,k,H,c)\geq\bigg{(}n-1-\frac{2(n-2)^{2}}{n^{2}-8}\bigg{)}\left(c+H^{2}\right)\geq\frac{n(n-1)}{n+2}\left(c+H^{2}\right),$
[…]
Next, it follows immediately from Lemma 2.5 (1) that the inequality (##\##2.1) is satisfied for $2\leq q\leq k$ at every point of $M$ and strict for $2\leq p<k$ at points with $c+H^{2}>0$. Then Lemma 2.1 implies that
$H_{q}(M;\mathbb{Z})=H_{n-q}(M;\mathbb{Z})=0\ \mbox{for each}\ 2\leq q\leq k-1.$
For $H_{k}(M;\mathbb{Z})$ and $H_{n-k}(M;\mathbb{Z})$, we divide the discussion into the following two cases.
Case (1): $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$. In this case, if $k=\left[\frac{n}{2}\right]$, then we know that $M$ is a homology sphere. And if $k<\left[\frac{n}{2}\right]$, then by the universal coefficient theorem, $H_{k}(M;\mathbb{Z})=0$ implies that $H^{k+1}(M;\mathbb{Z})$ is torsion-free, and thus $H_{n-k-1}(M;\mathbb{Z})$ is also torsion-free by the Poincaré duality theorem.
Case (2): $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold. It follows from Lemma 2.1 that for any point $x\in M$, there exists an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$ such that the equality holds in (##\##2.1) for $q=k$. Then by Lemma 2.5 (2), for any smooth orthonormal local frame $\left(\xi_{1},\cdots,\xi_{m}\right)$ for $NM$ over an open subset $U$ of $M$, the corresponding matrix
(3.2) $H_{\alpha}=\operatorname{diag}(\lambda_{\alpha}I_{k},\mu_{\alpha}I_{n-k})\ \mbox{for each}\ 1\leq\alpha\leq m,$
where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and $n-k$, respectively. It follows that the normal bundle $NM$ is flat. In addition, by Lemma 2.5 (2), we have
(3.3) $\operatorname{Ric}(e_{i},e_{j})=\delta_{ij}\alpha(n,k,H,c)\ \mbox{for any}\ 1\leq i,j\leq n.$
Since $M$ is connected and $n\geq 5$, by Schur’s lemma, we obtain that $M$ is an Einstein manifold and $\alpha(n,k,H,c)$ is constant. It follows that the mean curvature $H$ is also constant, and thus $c+H^{2}>0$ at every point of $M$. Note that the expressions of the principal normals
$\eta_{1}=\sum_{\alpha=1}^{m}\lambda_{\alpha}\xi_{\alpha}\ \mbox{and}\ \eta_{2}=\sum_{\alpha=1}^{m}\mu_{\alpha}\xi_{\alpha}$
are independent of the choice of the smooth orthonormal local frame $\left(\xi_{1},\cdots,\xi_{m}\right)$ for $NM$. Hence, there are two global normal vector fields, still denoted by $\eta_{1}$ and $\eta_{2}$, which are distinct principal normals of $f$ at every point.
In order to obtain the rigidity result by [9, Theorem 1], we need to prove that the mean curvature vector field $\mathcal{H}$ is parallel. For the case $H>0$, we can assume without loss of generality that $\xi_{1}=\mathcal{H}/{H}$. Then we have $\operatorname{tr}H_{1}=nH$ and $\operatorname{tr}H_{\alpha}=0$ for each $2\leq\alpha\leq m$. It follows from (3.2) that
(3.4) $k\lambda_{1}+(n-k)\mu_{1}=nH,$
(3.5) $k\lambda_{\alpha}+(n-k)\mu_{\alpha}=0$
for each $2\leq\alpha\leq m$. By (2.2) and (3.3), we have
(3.6) $\alpha(n,k,H,c)=(n-1)c+nH\lambda_{1}-\lambda_{1}^{2}-\sum_{\alpha=2}^{m}\lambda_{\alpha}^{2},$
(3.7) $\alpha(n,k,H,c)=(n-1)c+nH\mu_{1}-\mu_{1}^{2}-\sum_{\alpha=2}^{m}\mu_{\alpha}^{2}.$
Combining (3.4-3.7), one can verify by a direct calculation that
(3.8) $\lambda_{1}=H+(n-2k)\frac{\varphi(k)\left(c+H^{2}\right)}{kH},$
(3.9) $\mu_{1}=H-(n-2k)\frac{\varphi(k)\left(c+H^{2}\right)}{(n-k)H}$
are constant. It follows that $\sum\limits_{\alpha=2}^{m}\lambda_{\alpha}^{2},\sum\limits_{\alpha=2}^{m}\mu_{\alpha}^{2}$ are constant, and thus
$|\eta_{1}|^{2}=\lambda_{1}^{2}+\sum_{\alpha=2}^{m}\lambda_{\alpha}^{2},\ |\eta_{2}|^{2}=\mu_{1}^{2}+\sum_{\alpha=2}^{m}\mu_{\alpha}^{2},$
$\langle\eta_{1},\eta_{2}\rangle=\lambda_{1}\mu_{1}+\sum_{\alpha=2}^{m}\lambda_{\alpha}\mu_{\alpha}=\lambda_{1}\mu_{1}-\frac{k}{n-k}\sum_{\alpha=2}^{m}\lambda_{\alpha}^{2}$
are also constant. Since $N_{1}(x)=\operatorname{Span}\{\eta_{1}(x),\eta_{2}(x)\}$ for each $x\in M$ and the determinant of the Gram matrix $\left(\langle\eta_{i},\eta_{j}\rangle\right)_{2\times 2}$ is constant, we know that $f$ is $1$-regular and $\operatorname{rank}N_{1}=1$ or $2$. To prove the mean curvature vector field $\mathcal{H}$ is parallel, we divide the discussion into the following two cases.
Case (2-a): $\operatorname{rank}N_{1}=1$. In this case, it follows from (3.4-3.7) that $\lambda_{1}\mu_{1}\neq 0$ as $H>0$ and $\lambda_{\alpha}=\mu_{\alpha}=0$ for each $2\leq\alpha\leq m$. Then $\eta_{1}=\lambda_{1}\xi_{1}$ and $\eta_{2}=\mu_{1}\xi_{1}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and $n-k$, respectively. Therefore, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$. For every point of $M$, we can choose a smooth orthonormal local frame $\left(E_{1},\cdots,E_{n}\right)$ such that $E_{\eta_{1}}=\operatorname{Span}\{E_{1},\cdots,E_{k}\}$ and $E_{\eta_{2}}=\operatorname{Span}\{E_{k+1},\cdots,E_{n}\}$. It follows that
$\nabla_{E_{i}}^{\bot}\xi_{1}=\frac{1}{\lambda_{1}}\nabla_{E_{i}}^{\bot}\eta_{1}=0\ \mbox{for each}\ 1\leq i\leq k,$
$\nabla_{E_{j}}^{\bot}\xi_{1}=\frac{1}{\mu_{1}}\nabla_{E_{j}}^{\bot}\eta_{2}=0\ \mbox{for each}\ k+1\leq j\leq n,$
and thus $\mathcal{H}=H\xi_{1}$ is parallel.
Case (2-b): $\operatorname{rank}N_{1}=2$. In this case, we can assume additionally that $N_{1}\big{|}_{U}=\operatorname{Span}\{\xi_{1},\xi_{2}\}$. Then we have $\eta_{1}=\lambda_{1}\xi_{1}+\lambda_{2}\xi_{2}$ and $\eta_{2}=\mu_{1}\xi_{1}+\mu_{2}\xi_{2}$. It follows from (3.5-3.9) that $\lambda_{2},\mu_{2}$ are non-zero constants. Then $\eta_{1}$ and $\eta_{2}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and $n-k$, respectively. Similarly, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$. For every point of $M$, we can choose a smooth orthonormal local frame $\left(E_{1},\cdots,E_{n}\right)$ such that $E_{\eta_{1}}=\operatorname{Span}\{E_{1},\cdots,E_{k}\}$ and $E_{\eta_{2}}=\operatorname{Span}\{E_{k+1},\cdots,E_{n}\}$. Note that (3.2) implies that
$k\eta_{1}+(n-k)\eta_{2}=n\mathcal{H},$
and thus we have
$\operatorname{II}(E_{1},E_{1})=\eta_{1}=\frac{n}{k}\mathcal{H}-\frac{n-k}{k}\eta_{2}=\frac{1}{k}\sum_{i=1}^{n}\operatorname{II}(E_{i},E_{i})-\frac{n-k}{k}\operatorname{II}(E_{n},E_{n}),$
i.e.,
$(k-1)\operatorname{II}(E_{1},E_{1})-\sum_{i=2}^{n-1}\operatorname{II}(E_{i},E_{i})+(n-k-1)\operatorname{II}(E_{n},E_{n})=0.$
It follows from Lemma 2.4 that $N_{1}$ is parallel, i.e., $\nabla_{X}^{\bot}\xi\subset N_{1}$ for any $X\in TM$ and any smooth section $\xi$ of $N_{1}$. By $\langle\xi_{1},\xi_{1}\rangle=\langle\xi_{2},\xi_{2}\rangle=1$, we obtain
(3.10) $\langle\nabla_{E_{i}}^{\bot}\xi_{1},\xi_{1}\rangle=\langle\nabla_{E_{i}}^{\bot}\xi_{2},\xi_{2}\rangle=0\ \mbox{for each}\ 1\leq i\leq n.$
Since $\eta_{1},\eta_{2}$ are Dupin principal normal vector fields, we have
(3.11) $\nabla_{E_{i}}^{\bot}\eta_{1}=\lambda_{1}\nabla_{E_{i}}^{\bot}\xi_{1}+\lambda_{2}\nabla_{E_{i}}^{\bot}\xi_{2}=0\ \mbox{for each}\ 1\leq i\leq k,$
(3.12) $\nabla_{E_{j}}^{\bot}\eta_{2}=\mu_{1}\nabla_{E_{j}}^{\bot}\xi_{1}+\mu_{2}\nabla_{E_{j}}^{\bot}\xi_{2}=0\ \mbox{for each}\ k+1\leq j\leq n.$
Combining (3.10-3.12) and the fact that $N_{1}$ is parallel, we have $\nabla_{E_{i}}^{\bot}\xi_{1}=0$ for each $1\leq i\leq n$, and thus $\mathcal{H}=H\xi_{1}$ is parallel.
To sum up, we have already proved that $M$ is an Einstein submanifold with flat normal bundle and parallel mean curvature vector field. Then by [9, Theorem 1], there exist an isometric immersion
$g:M^{n}\rightarrow S^{n_{1}}\left(r_{1}\right)\times\cdots\times S^{n_{l}}\left(r_{l}\right)\subset S^{n+l-1}\big{(}1/\sqrt{\overline{c}}\,\big{)}$
and a totally umbilical embedding
$u:S^{n+l-1}\big{(}1/\sqrt{\overline{c}}\,\big{)}\rightarrow F^{n+m}(c)$
such that $f=u\circ g$ and $\alpha(n,k,H,c)=(n-l)\overline{c}$ for an positive integer $l$, where $r_{i}^{2}=\frac{n_{i}-1}{\alpha(n,k,H,c)}$, $2\leq n_{i}\leq n$ and $\sum_{i=1}^{l}n_{i}=n$. Since $u$ is totally umbilical, it is easy to verify that $H^{2}=H_{g}^{2}+H_{u}^{2}$ and $\overline{c}=c+H_{u}^{2}$, where $H_{g}$ and $H_{u}$ denote the mean curvature of $g$ and $u$, respectively. It follows that $\alpha(n,k,H,c)=(n-l)\left(c+H_{u}^{2}\right)\leq(n-l)\left(c+H^{2}\right)$. Then by Lemma 2.3, we have $n-3<n-l$, i.e., $l=1$ or $2$. Recall that $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold in Case (2). Hence, we must have $l=2$ and $M$ is isometric to the Einstein hypersurface $S^{k}\left(\sqrt{\frac{k-1}{n-2}}\,r\right)\times S^{n-k}\left(\sqrt{\frac{n-k-1}{n-2}}\,r\right)$ in the totally umbilical sphere $S^{n+1}\left(r\right)\subset F^{n+m}(c)$, where $r=\sqrt{\frac{n-2}{\alpha(n,k,H,c)}}$. Finally, one can verify by a direct calculation that the Ricci curvature and the mean curvature of this Einstein hypersurface are exactly
$\operatorname{Ric}=\alpha(n,k,H,c)=\frac{n-2}{r^{2}}$
and
$H_{g}=\sqrt{H^{2}+c-\overline{c}}=\sqrt{\left(\varphi(k)-\frac{1}{n-2}\right)\left(c+H^{2}\right)}=\frac{n-2k}{rn\sqrt{(k-1)(n-k-1)}}.$
3.2. Proof of Corollary 1.3
Proof.
Note that Lemma 2.3 implies that $(n-2)\left(c+H^{2}\right)\geq\alpha(n,\left[\frac{n}{2}\right],H,c)$ and the equality holds if and only if $n=2k$. Then by Theorem 1.2 with $k=\left[\frac{n}{2}\right]$, $M$ is either a simply connected homology sphere or isometric to the minimal Clifford hypersurface $S^{k}\left(r/\sqrt{2}\right)\times S^{k}\left(r/\sqrt{2}\right)$ in the totally umbilical sphere $S^{n+1}\left(r\right)$, where $n=2k$ and $r=\frac{1}{\sqrt{c+H^{2}}}$. It follows from the Hurewicz theorem, the Whitehead theorem and the generalized Poincaré conjecture that every simply connected homology sphere is a topological sphere. ∎
3.3. Proof of Corollary 1.4
Proof.
Note that for $n=2k+1$, we have
$\alpha(n,k,H,c)=\left(n-2-\frac{4}{n^{3}-2n^{2}-n-2}\right)\left(c+H^{2}\right).$
Then the proof is completed by the same argument as in the proof of Corollary 1.3. ∎
Acknowledgments.
The authors would like to thank Professor Marcos Dajczer, Professor Theodoros Vlachos and Professor Rodrigo Montes for their interest and comments.
Data Availibility.
This manuscript has no associated data.
Conflict of interest.
On behalf of all authors, the corresponding author states that there is no conflict of interest.

### v3

@@
 the corresponding author.
-J. Q. Ge is partially supported by NSFC (No. 12171037) and the Fundamental Research Funds for the Central Universities.
-Y. Zhou is partially supported by NSFC (No. 12171037, 12271040), China Postdoctoral Science Foundation (No. BX20230018) and National Key R&\&&D Program of China 2020YFA0712800.
+J. Q. Ge is partially supported by NSFC (No. 12171037, 12571049) and the Fundamental Research Funds for the Central Universities.
+Y. Zhou is partially supported by NSFC (No. 12171037, 12271040), China Postdoctoral Science Foundation (No. BX20230018) and National Key R&\&D Program of China 2020YFA0712800.
 1. Introduction
@@
 In the same paper, Xu-Gu raised a conjecture, indicating that the above generalization should also hold without the assumption of parallel mean curvature vector field.
-Recently, Dajczer-Vlachos [3] affirmed Xu-Gu’s conjecture for the case $n\geq 4$, expect for the part of diffeomorphic to spheres.
+Recently, Dajczer-Vlachos [3] affirmed Xu-Gu’s conjecture for the case $n\geq 4$, except for the part of diffeomorphic to spheres.
 The same philosophy was used in a series of rigidity results [1, 2, 4] for submanifolds with pinched Ricci curvature in different settings.
@@
 Theorem 1.1.
-(((([12, Theorem 2])))) Let $M^{n}$ be an $n(\geq 4)$-dimensional compact submanifold of the unit sphere $\mathbf{S}^{n+m}$. Assume that the Ricci curvature satisfies
+(([12, Theorem 2])) Let $M^{n}$ be an $n(\geq 4)$-dimensional compact submanifold of the unit sphere $\mathbf{S}^{n+m}$. Assume that the Ricci curvature satisfies
 (1.1) $\operatorname{Ric}>b(n,k,H):=\frac{n(k-1)}{k}+\frac{n(k-1)H}{2k^{2}}\left(nH+\sqrt{n^{2}H^{2}+4k(n-k)}\right),$
@@
 Theorem 1.2.
-Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 5$, $c\geq 0$, be an isometric immersion of a compact, connected Riamannian manifold. If the Ricci curvature of $M$ satisfies
+Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 5$, $c\geq 0$, be an isometric immersion of a compact, connected Riemannian manifold. If the Ricci curvature of $M$ satisfies
 (*) $\operatorname{Ric}\geq\alpha(n,k,H,c)=\bigg(n-1-\dfrac{k(n-k)(n-2)}{k(n-2k)+n(k-1)(n-k)}\bigg)\left(c+H^{2}\right)$
@@
 Next, we will present several lemmas used in the proof of Theorem 1.2 and Corollary 1.3.
-The following Lemma 2.1 was first proved by Lawson-Simons [8] for the case $c>0$, and then by Xin [13] for the case $c=0$ when the inequality (##\##2.1) is strict on the whole submanifold.
-Elworthy-Rosenberg [7] observed that the result still holds by only requiring the inequality (##\##2.1) is strict at one point.
+The following Lemma 2.1 was first proved by Lawson-Simons [8] for the case $c>0$, and then by Xin [13] for the case $c=0$ when the inequality (#\#2.1) is strict on the whole submanifold.
+Elworthy-Rosenberg [7] observed that the result still holds by only requiring the inequality (#\#2.1) is strict at one point.
 Lemma 2.1.
-(((([7, 8, 13])))). Let $f:M^{n}\rightarrow F^{n+m}(c)$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that
-(##\##) $\Theta_{q}:=\sum_{i=1}^{q}\sum_{j=q+1}^{n}\left(2|\operatorname{II}(e_{i},e_{j})|^{2}-\langle\operatorname{II}(e_{i},e_{i}),\operatorname{II}(e_{j},e_{j})\rangle\right)\leq q(n-q)c$
-holds for any point $x\in M$ and any orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$, where $q$ is an integer such that $11\leq q\leq n-$ 1.
-If there exists a point of $M$ such that the inequality ((((#italic-#\##2.1)))) is strict, then $H_{q}(M;\mathbb{Z})=H_{n-q}(M;\mathbb{Z})=0$.
+(([7, 8, 13])). Let $f:M^{n}\rightarrow F^{n+m}(c)$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that
+(#\#) $\Theta_{q}:=\sum_{i=1}^{q}\sum_{j=q+1}^{n}\left(2|\operatorname{II}(e_{i},e_{j})|^{2}-\langle\operatorname{II}(e_{i},e_{i}),\operatorname{II}(e_{j},e_{j})\rangle\right)\leq q(n-q)c$
+holds for any point $x\in M$ and any orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$, where $q$ is an integer such that $1\leq q\leq n-1$.
+If there exists a point of $M$ such that the inequality ((#\#2.1)) is strict, then $H_{q}(M;\mathbb{Z})=H_{n-q}(M;\mathbb{Z})=0$.
 Using Lemma 2.1, Dajczer-Vlachos [3] improved [14, Lemma 2.2] to the following Lemma 2.2.
 Lemma 2.2.
-(((([3, Lemma 4])))) Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 4$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that the Ricci curvature of $M$ satisfies
+(([3, Lemma 4])) Let $f:M^{n}\rightarrow F^{n+m}(c)$, $n\geq 4$, $c\geq 0$, be an isometric immersion of a compact manifold. Suppose that the Ricci curvature of $M$ satisfies
 (2.4) $\operatorname{Ric}\geq\frac{n(n-1)}{n+2}\left(c+H^{2}\right).$
@@
 Lemma 2.4.
-(((([10, Lemma 5])))) Let $f:M^{n}\rightarrow F^{n+m}(c)$ be a $1$-regular immersion.
+(([10, Lemma 5])) Let $f:M^{n}\rightarrow F^{n+m}(c)$ be a $1$-regular isometric immersion.
 If for every point of $M$, there exists an orthonormal frame $\{E_{1},\cdots,E_{n}\}$ such that for each $1\leq i\leq n$, $\operatorname{II}(E_{i},E_{j})=0$ for any $j\neq i$ and $\operatorname{II}(E_{i},E_{i})=\sum_{j\neq i}a_{j}\operatorname{II}(E_{j},E_{j})$ for some real numbers $a_{j}$’s, then the first normal bundle $N_{1}$ is parallel.
@@
   • (1)
-    The inequality ((((#italic-#\##2.1)))) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$.
-Moreover, if the inequality ((((*2.5)))) is strict at $x$, then the inequality ((((#italic-#\##2.1)))) is strict for $q=k$ at $x$.
+    The inequality ((#\#2.1)) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$.
+Moreover, if the inequality ((*2.5)) is strict at $x$, then the inequality ((#\#2.1)) is strict for $q=k$ at $x$.
   • (2)
-    Suppose that the equality holds in ((((#italic-#\##2.1)))) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$.
+    Suppose that the equality holds in ((#\#2.1)) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$.
 Then for any orthonormal basis $\{\xi_{1},\cdots,\xi_{m}\}$ of $N_{x}M$, the corresponding matrix
     $H_{\alpha}=\operatorname{diag}(\lambda_{\alpha}I_{k},\mu_{\alpha}I_{n-k})\ \mbox{for each}\ 1\leq\alpha\leq m,$
-    where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and $n-k$, respectively.
+    where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and nkn-k, respectively.
 Furthermore, if $c+H^{2}>0$ at $x$, then $\eta_{1}=\sum_{\alpha=1}^{m}\lambda_{\alpha}\xi_{\alpha}$ and $\eta_{2}=\sum_{\alpha=1}^{m}\mu_{\alpha}\xi_{\alpha}$ are distinct principal normals of $f$ at $x$ such that $E_{\eta_{1}}(x)=\operatorname{Span}\{e_{1},\cdots,e_{k}\}$ and $E_{\eta_{2}}(x)=\operatorname{Span}\{e_{k+1},\cdots,e_{n}\}$.
@@
 Note that the last inequality in (2.11) is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$.
-Hence the inequality (##\##2.1) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$.
-If in addition the inequality (*2.5) is strict at $x$, then the second inequality in (2.11) is strict, and thus the inequality (##\##2.1) is strict for $q=k$ at $x$.
+Hence the inequality (#\#2.1) is satisfied for $2\leq q\leq k$ at $x$ and is strict for $2\leq q<k$ if $c+H^{2}>0$ at $x$.
+If in addition the inequality (*2.5) is strict at $x$, then the second inequality in (2.11) is strict, and thus the inequality (#\#2.1) is strict for $q=k$ at $x$.
 Next, we prove $(2)$.
-Assume that the equality holds in (##\##2.1) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$.
+Assume that the equality holds in (#\#2.1) with $q=k$ for an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$.
 Then all inequalities in (2.9-2.11) become equalities.
@@
 […]
-Next, it follows immediately from Lemma 2.5 (1) that the inequality (##\##2.1) is satisfied for $2\leq q\leq k$ at every point of $M$ and strict for $2\leq p<k$ at points with $c+H^{2}>0$.
+Next, it follows immediately from Lemma 2.5 (1) that the inequality (#\#2.1) is satisfied for $2\leq q\leq k$ at every point of $M$ and strict for $2\leq p<k$ at points with $c+H^{2}>0$.
 Then Lemma 2.1 implies that
@@
 Case (2): $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold.
-It follows from Lemma 2.1 that for any point $x\in M$, there exists an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$ such that the equality holds in (##\##2.1) for $q=k$.
+It follows from Lemma 2.1 that for any point $x\in M$, there exists an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ of $T_{x}M$ such that the equality holds in (#\#2.1) for $q=k$.
 Then by Lemma 2.5 (2), for any smooth orthonormal local frame $\left(\xi_{1},\cdots,\xi_{m}\right)$ for $NM$ over an open subset $U$ of $M$, the corresponding matrix
 (3.2) $H_{\alpha}=\operatorname{diag}(\lambda_{\alpha}I_{k},\mu_{\alpha}I_{n-k})\ \mbox{for each}\ 1\leq\alpha\leq m,$
-where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and $n-k$, respectively.
+where $\lambda_{\alpha}$ and $\mu_{\alpha}$ are principal curvatures with respect to $\xi_{\alpha}$ of multiplicities $k$ and nkn-k, respectively.
 It follows that the normal bundle $NM$ is flat.
@@
 In this case, it follows from (3.4-3.7) that $\lambda_{1}\mu_{1}\neq 0$ as $H>0$ and $\lambda_{\alpha}=\mu_{\alpha}=0$ for each $2\leq\alpha\leq m$.
-Then $\eta_{1}=\lambda_{1}\xi_{1}$ and $\eta_{2}=\mu_{1}\xi_{1}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and $n-k$, respectively.
+Then $\eta_{1}=\lambda_{1}\xi_{1}$ and $\eta_{2}=\mu_{1}\xi_{1}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and nkn-k, respectively.
 Therefore, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$.
@@
 It follows from (3.5-3.9) that $\lambda_{2},\mu_{2}$ are non-zero constants.
-Then $\eta_{1}$ and $\eta_{2}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and $n-k$, respectively.
+Then $\eta_{1}$ and $\eta_{2}$ are smooth normal vector fields and hence Dupin principal normal vector fields with multiplicities $k$ and nkn-k, respectively.
 Similarly, $E_{\eta_{1}}$ and $E_{\eta_{2}}$ are smooth distributions such that $TM=E_{\eta_{1}}\oplus E_{\eta_{2}}$.
@@
 $(k-1)\operatorname{II}(E_{1},E_{1})-\sum_{i=2}^{n-1}\operatorname{II}(E_{i},E_{i})+(n-k-1)\operatorname{II}(E_{n},E_{n})=0.$
-It follows from Lemma 2.4 that $N_{1}$ is parallel, i.e., $\nabla_{X}^{\bot}\xi\subset N_{1}$ for any $X\in TM$ and any smooth section $\xi$ of $N_{1}$.
+It follows from Lemma 2.4 that $N_{1}$ is parallel, i.e., $\nabla_{X}^{\bot}\xi\in\Gamma(N_{1})$ for any $X\in TM$ and any smooth section $\xi\in\Gamma(N_{1})$.
 By $\langle\xi_{1},\xi_{1}\rangle=\langle\xi_{2},\xi_{2}\rangle=1$, we obtain
@@
 It follows that $\alpha(n,k,H,c)=(n-l)\left(c+H_{u}^{2}\right)\leq(n-l)\left(c+H^{2}\right)$.
-Then by Lemma 2.3, we have $n-3<n-l$, i.e., $l=1$ or $2$.
+Then by Lemma 2.3, we have n3<nln-3<n-l, i.e., $l=1$ or $2$.
 Recall that $H_{k}(M;\mathbb{Z})=H_{n-k}(M;\mathbb{Z})=0$ does not hold in Case (2).
@@
 The authors would like to thank Professor Marcos Dajczer, Professor Theodoros Vlachos and Professor Rodrigo Montes for their interest and comments.
-Data Availibility.
-This manuscript has no associated data.
-Conflict of interest.
-On behalf of all authors, the corresponding author states that there is no conflict of interest.